OpenStudy (anonymous):

Find the missing value such that the given binomial is a factor of the given polynomial. x^2-4x-?, X+6 please show the work

OpenStudy (anonymous):

yes you want \[x^2-4x+?=(x+6)(something)\]

OpenStudy (anonymous):

when you substitute x = -6 you will have to get 0 as an answer because the first term will be 0. so substitute x = -6 on the left hand side, set = 0 and solve

OpenStudy (anonymous):

ahhh i see

OpenStudy (anonymous):

\[(-6)^2-4\times (-6)+?=0\] \[36+24+?=0\] \[60+?=0\] \[?=-60\]

OpenStudy (anonymous):

so it is \[x^2-4x-60=(x+6)(x-10)\]

OpenStudy (anonymous):

k?

OpenStudy (anonymous):

got it thanks!

OpenStudy (anonymous):

welcome